OpenStudy (anonymous):

Nafeesa Graphed a line with slope 5 and y-intercept of -2. a) What is the equation of her line? b) Find the value of X when y=0.

OpenStudy (tkhunny):

Please show your best efforts. You got one for free. Now, it's your turn.

OpenStudy (anonymous):

I am trying

OpenStudy (tkhunny):

You should have a Slope-Intercept form of a line. Do you have it?

OpenStudy (wolf1728):

The equation would be b = y -mx we know the slope = 5 and so the equation becomes b = y -5x Can you work it out from there?

OpenStudy (wolf1728):

The Y-intercept is -2, so we have a point to work with - when x=0 y = -2 so that point is (0,-2) Plugging the point into the equation we have b = y -mx b = -2 -5*0 b = -2 -2 = y -5x y = 5x -2